# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v2

@@ 4.2.1. Convergence analysis of the energy
We notice that the energy convergence consists of a deterministic power iteration part and a fluctuating part that is rougly $a_{0}\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}=\left|c_{\perp}/c_{0}\right|\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}$ (when $a_{0}$ is small). Here $\|H-E_{0}I\|_{2}$ resembles the energy scale, and $\tan\theta_{k}^{(n)}$ is a random variable that captures the deviation of a random walker from the ground state. While currently, we do not have a characterization of the variance of $\tan\theta_{k}^{(n)}$, if we assume that it has a bounded variance of $O(1)$, when $a_{0}$ gets small, i.e. the trial function gets closer to the ground-state wavefunction, the fluctuation term can be made small. In an ideal case where $a_{0}=0$ so that the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is no fluctuation in the energy estimation. As a matter of fact, we have
$\frac{\braket{\Psi_{0},H\Psi}}{\braket{\Psi_{0},\Psi}}=E_{0},$
for any wavefunction $\Psi$.
4.2.2. Convergence analysis of the walkers
Now, we analyze the convergence of a particular walker $\Phi_{k}^{(n)}$ to the ground state, by keeping track of the angle between $\Phi_{k}^{(n)}$ and the ground state $\Psi_{0}$. As shown later in Theorem 6, unlike the results in energy convergence, even if the trial function is exactly the ground-state wavefunction, there is still an error term that does not vanish. This implies that a single walker may not be able to converge to somewhere near the ground state, even with a good trial wavefunction.
Theorem 6.
Then,
$\tan\theta_{k}^{(n+1)}(\vec{x})\leq e^{-\Delta\tau(E_{1}-E_{0})}\tan\theta_{k}^{(n)}+e^{\Delta\tau E_{0}}\left(a_{0}\tan\theta_{k}^{(n)}+1\right)O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)\\ \quad+a_{0}e^{2\Delta\tau E_{0}}O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)^{2},$
Proof.
@@ 4.2.2. Convergence analysis of the walkers
As we can see, even when $a_{0}=0$, i.e. the trial wavefunction is exactly the ground-state wavefunction, there is still an error term $O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$. Thus, the walker is not guaranteed to get closer to the ground state. In fact, generally, one needs exponential number of walkers to get $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}<1$. This is due to the fact that generally, one needs to sample an exponential number of $\widetilde{B}_{k}^{(n)}(\vec{x})$ to approximate $\exp(-\Delta\tau H)$.
5. Numerical Results
In this section, we apply our proposed algorithm to both 1D and 2D Transverse-Field Ising (TFI) systems. For more details about TFI system, we refer readers to Appendix A. For the 1D system, we consider a periodic boundary condition, while for the 2D system, we use an open boundary condition.
In all of our experiments, the step size of the imaginary-time evolution is set to be $\Delta\tau=0.01$. We set the initial trial wavefunction to be the fully disordered state
(42) $\Psi_{\mathrm{tr}}=\frac{1}{\sqrt{2^{d}}}\cdot\left(1,1\right)^{T}\otimes\cdots\otimes\left(1,1\right)^{T},$
which can also be interpreted as the eigenvector corresponding to the smallest eigenvalue of $H_{1}$ (The first term of the Hamiltonian in (43)). We implement TT-sketching to update the trial wavefunction every $50$ steps. At the $2000$-th step, we stop updating the trial wavefunction and keep using the latest one. We set the number of walkers to scale linearly with the number of spins. For systems with $16,32,64$ and $96$ spins, the numbers of walkers are set to be $2000,4000,8000$ and $12000$, respectively.
In the TT-sketching procedure, we take $\delta=0.1$ in (15) for the sketching functions. For systems with $16$ or $32$ spins, the rank of the sketching functions is set to be $60$. For systems with $64$ or $96$ spins, we set the rank to be $150$. The target rank of the sketched TT is set to be $r_{1}=r_{d-1}=2$, and $r_{2}=\cdots=r_{d-2}=4$ in all experiments.
The 1D transverse-field Ising model with periodic boundary undergoes a quantum phase transition [45] at $g=1$ (see (43)), which is demonstrated in Figure 3 using the 1D system with $16$ spins as an example. To be more specific, We run our algorithm on the 1D system with $16$ spins with different transverse fields, and measure the ground-state energy $E_{0}(g)$. We also calculate the energy of the transverse field by
where we pick $\Delta g=0.01$ in practice.
Figure 3. Results for $16$ spins in 1D with different transverse-field strength $g$. Left: ground-state energy; right: the energy of the transverse field, where the critical point $g=1$ is illustrated with the black dashed line.
Since the quantum phase transition at the critical point $g=1$ makes it the most difficult case to deal with, we also compare cp-AFQMC with re-anchoring and the vanilla cp-AFQMC on larger systems, including 1D system with $d=32,64$ and $96$ spins at $g=1$. The ground-state energy of 1D periodic system can be calculated analytically [46], and the ground-state wavefunction can be well approximated by density matrix renormalization group (DMRG) [47, 48].
In addition, we also test a quasi-1D system, where the spins are placed on a $4\times 16$ lattice. We incorporate periodic boundary on the $4$-site axis, and open boundary on the $16$-site axis. Thus, it appears like a cylinder. The transverse field is also set to be $g=1$. Despite that the spins are arranged on a 2D lattice, the ground-state energy and ground-state wavefunction can still be well-approximated by DMRG, due to the specific type of boundary conditions. Thus, we still use DMRG result as a reference.
As illustrated in Figure 4, for each system, the energy obtained by vanilla cp-AFQMC is of the order $O(10^{-3})$, while cp-AFQMC with re-anchoring can achieve the accuracy of order $O(10^{-5})$ or $O(10^{-6})$. The relative error of the energy is shown in Table 1. Besides, the overlap between our optimized trial wavefunction and the ground-state wavefunction is shown in Figure 5. For 1D systems, the optimized trial wavefunction in the form of a rank-$4$ TT can approximate the ground-state wavefunction quite well. For systems with $32$, $64$ and $96$ spins, the overlap can achieve around $0.9$, $0.8$, and $0.7$ respectively. For the $4\times 16$ cylinder system, the trial wavefunction may converge to somewhere not quite close to the ground-state wavefunction. However, we can still get pretty accurate ground-state energy using the optimized trial wavefunction.
Figure 4. Comparison between the energy convergence of cp-AFQMC and cp-AFQMC with re-anchoring at $g=1$. Top: cp-AFQMC; bottom: cp-AFQMC with re-anchoring. From left to right: $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
Figure 5. The overlap between our optimized trial wavefunction and the ground-state wavefunction. From left to right: $32$ spins in 1D; $64$ spins in 1D; $96$ spins in 1D; $4\times 16$ spins in 2D (cylinder).
cp-AFQMC cp-AFQMC with re-anchoring
$32$ spins 1D $(+1.35\pm 0.31)\times 10^{-3}$ $(+0.44\pm 2.43)\times 10^{-5}$
$64$ spins 1D $(+1.88\pm 0.37)\times 10^{-3}$ $(+0.77\pm 1.07)\times 10^{-5}$
$96$ spins 1D $(+2.49\pm 0.33)\times 10^{-3}$ $(-4.95\pm 8.94)\times 10^{-6}$
$4\times 16$ spins 2D (cylinder) $(+4.18\pm 1.42)\times 10^{-3}$ $(-0.66\pm 1.96)\times 10^{-6}$
Table 1. Relative error of the energy.
From the results in Table 1, we can tell that there is an obvious bias in the estimation of the ground-state energy by vanilla cp-AFQMC. As discussed in detail in Appendix B, the bias originates from Trotter error and population control.
Finally, we test our algorithm on the $11\times 1$ 11 lattice with open boundary at the quantum critical point $g\approx 3.044$ [49]. In our experiments, we take $g=3$. The results are shown in Figure 6. For such a 2D system, there is no analytic solution for the ground-state energy anymore. A benchmark for the energy density given by DMRG is $-3.17210\pm 1\times 10^{-5}$ [50], while the energy density estimated by our method is $-3.17212\pm 4\times 10^{-5}$.
Figure 6. Energy convergence of cp-AFQMC with re-anchoring for $11\times 1$ 11 lattice.
6. Conclusion
In this paper, we proposed a framework for calculating the ground-state energy of quantum many-body systems. Our algorithm improved cp-AFQMC by periodically updating the trial wavefunction using TT-sketching. As indicated by numerical results, our algorithm is accurate for the transverse-field Ising model. Furthermore, we provided a convergence analysis for cp-AFQMC for the first time. A potential work in the future is to apply this framework to electronic systems.
Appendix A Transverse-Field Ising Model
We take the transverse-field Ising (TFI) model [51, 52] as an example. We follow the procedure of the CPMC algorithm on Hubbard model [41] to solve for TFI model. The Hamiltonian of TFI model is given by
For such a Hamiltonian, we use the Suzuki-Trotter approximation [54, 55] to approximate $e^{-\Delta\tau H}$:
@@ Appendix A Transverse-Field Ising Model
      Update $\phi_{k}^{(n)\prime}$ to $\phi_{k}^{(n)\prime\prime}$ by applying $B_{H_{2}}$ to $\phi_{k}^{(n)\prime}$ stochastically using the decomposition in (46). Update weights $w_{k}^{(n)}$ to $w_{k}^{(n+1)}$. The details are provided in Section A.1.
@@
  • Periodically compute energy as in Section 2.1.1.
  • Periodically do population control as in Section 2.1.2.
@@
We remark that all of the entries in both the matrices $B_{H_{1}/2}$ and $B_{H_{2}}$ are non-negative. Thus, if we apply $B_{H_{1}/2}B_{H_{2}}B_{H_{1}/2}$ to a wavefunction with all the entries being positive, it returns another all-positive-entry function.
A.1. Details for applying $B_{H_{2}}$
@@
Appendix B Systematic bias of vanilla cp-AFQMC on TFI system
In this section, we study the origin of the bias in this section, using the TFI system with $16$ spins in 1D as an example.
The good news is that for TFI system, with a properly chosen trial wavefunction $\Psi_{\mathrm{tr}}$, the cp-AFQMC is free of sign problem. In fact, as shown in Appendix A, the operator $B(\vec{x})$ in (4) has the property that if it acts on a wavefunction $\phi$ with all entries being positive, the result $B(\vec{x})\phi$ still has all positive entries. Therefore, by choosing a trial function with all positive entries, the walkers will remain such a form, thus it always holds $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}>0$. The trial wavefunction (42) satisfies the condition, thus there will be no sign problem if one runs cp-AFQMC with this fixed trial function.
Without sign problem, the constrained-path approximation will not lead to any systematic bias by zeroing some of the walkers. However, there are other issues that will bring a bias to the energy estimation. One of them is the Trotter error as shown in (44). At each imaginary-time step, the Trotter error is $O(\Delta\tau^{3})$. For a fixed imaginary-time, one needs to run $O(1/\Delta\tau)$ steps of cp-AFQMC. Thus, the total error in terms of the energy estimation accumulates to $O(\Delta\tau^{2})$.
Ideally, the Trotter error will vanish as the imaginary-time step $\Delta\tau$ goes to 00. However, this is not always the case as shown in Figure 7(a). With $2000$ walkers, the estimated energy indeed converges to the ground-state energy as $\Delta\tau$ goes to 00. If we use $200$ walkers, however, there is still a quite obvious bias even when $\Delta\tau$ becomes 00. Moreover, with fewer walkers, the bias is consistently larger with any imaginary-time step $\Delta\tau$.
@@ Appendix B Systematic bias of vanilla cp-AFQMC on TFI system
To further study how the error depends on the number of walkers, we fix the imaginary-time step to be $\Delta\tau=0.01$, and run cp-AFQMC with different population size. As shown in Figure 7(b), the error is roughly inversely proportional to the population size. The bias is introduced by the population control procedure summarized in Algorithm 5. To be more specific, if we look at each single walker, it breaks into a random number of copies, with the average number of copies matching its weight. However, to maintain a constant population size, the number of copies must be adjusted based on the total, which introduces dependency among walkers. This dependency leads to a bias in the energy estimation.

### v4

@@ Appendix B Details on Convergence Analysis of cp-AFQMC
Proof of Theorem 6.
@@
Appendix C cp-AFQMC on Transverse-Field Ising Model
In this section, we follow the procedure of the CPMC algorithm on Hubbard model [57] to solve for TFI model. For the Hamiltonian (27), we use the Suzuki-Trotter approximation [58, 59] to approximate $e^{-\Delta\tau H}$:
@@
      Update $\phi_{k}^{(n)\prime}$ to $\phi_{k}^{(n)\prime\prime}$ by applying $B_{H_{2}}$ to $\phi_{k}^{(n)\prime}$ stochastically using the decomposition in (49). Update weights $w_{k}^{(n)}$ to $w_{k}^{(n+1)}$. The details are provided in Section C.1.
@@
  • Periodically estimate the ground-state energy.
  • Periodically do population control.
@@
We remark that all of the entries in both the matrices $B_{H_{1}/2}$ and $B_{H_{2}}$ are non-negative. Thus, if we apply $B_{H_{1}/2}B_{H_{2}}B_{H_{1}/2}$ to a wavefunction with all the entries being positive, it returns another all-positive-entry function. Therefore, by choosing a trial function with all positive entries, the walkers will remain such a form, thus it always holds $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}>0$. The trial wavefunction (28) satisfies the condition, thus there will be no sign problem if one runs cp-AFQMC with this fixed trial function.
C.1. Details for applying $B_{H_{2}}$
@@
Appendix D Systematic Error of Vanilla cp-AFQMC on TFI System
In this section, we study the origin of the bias in cp-AFQMC, using the TFI system with $16$ spins in 1D as an example. As mentioned earlier in Section 5, the TFI system is free of sign problem with a properly chosen trial wavefunction, but there is still a systematic error caused by the fixed imaginary-time step $\Delta\tau$ and the finite population size $N$.
At each iteration step, a Trotter error $O(\Delta\tau^{3})$ is introduced, as shown in (47). For a fixed imaginary-time, one needs to run $O(1/\Delta\tau)$ steps of random walk. Thus, the total error in terms of the energy estimation accumulates to $O(\Delta\tau^{2})$. Ideally, the Trotter error will vanish as the imaginary-time step $\Delta\tau$ goes to 00. However, this is not always the case as shown in Figure 14(a). With $2000$ walkers, the estimated energy indeed converges to the ground-state energy as $\Delta\tau$ goes to 00. If we use $200$ walkers, however, there is still a quite obvious bias even when $\Delta\tau$ becomes 00. Moreover, with fewer walkers, the bias is consistently larger with any imaginary-time step $\Delta\tau$.
@@ Appendix D Systematic Error of Vanilla cp-AFQMC on TFI System
To further study how the error depends on the number of walkers, we fix the imaginary-time step to $\Delta\tau=0.01$, and run cp-AFQMC with different population size. As shown in Figure 14(b), the error is roughly inversely proportional to the population size. The bias is introduced by the population control procedure summarized in Algorithm 5. To be more specific, if we look at each single walker, it breaks into a random number of copies, with the average number of copies matching its weight. However, to maintain a constant population size, the number of copies must be adjusted based on the total, which introduces dependency among walkers. This dependency leads to a bias in the energy estimation.